OpenStudy (anonymous):

Find the slope of the line that passes through the pair of points. A(8, 1), B(-1, -7)

OpenStudy (anonymous):

use the slope formula : slope = (y2 - y1) / (x2 - x1) (8,1) x1 = 8 and y1 = 1 (-1,-7) x2 = -1 and y2 = -7 now we sub slope = (-7-1) / (-1-8) slope = -8/-9 slope = 8/9 any questions ?